OpenStudy (anonymous):

c*9c need help will medal and fan

OpenStudy (anonymous):

c * 9c

OpenStudy (anonymous):

@angelg139 help plz

Directrix (directrix):

c * 9c means c times 9 times c

OpenStudy (anonymous):

so it the same

Directrix (directrix):

c times c = what?

OpenStudy (anonymous):

9

Directrix (directrix):

I need to ask what the instructions are for this problem.

Directrix (directrix):

c times c is not 9 c*c = c^2

OpenStudy (anonymous):

write the product in its simplest form. c * 9c

Directrix (directrix):

And, then 9 times that

OpenStudy (anonymous):

81

OpenStudy (anonymous):

im so losted

Directrix (directrix):

9 * c^2 = 9 c²

OpenStudy (anonymous):

thx

Directrix (directrix):

You are welcome.

OpenStudy (anonymous):

so when it is like that you add

Directrix (directrix):

Yes c^1 * c^1 = c^(1 + 1) = c²

OpenStudy (anonymous):

6c * (-7c^6) how do you do this one

Directrix (directrix):

And, then multiply that by 9 to get: 9c²

Directrix (directrix):

Give it a try.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

we add 6 to both sides

Directrix (directrix):

This is not an equation so there are no sides.

OpenStudy (anonymous):

ohhh I feel dum

Directrix (directrix):

6c * (-7c^6) = 6 * (-7) * c * c^6 =

Directrix (directrix):

What is 6 times negative 7 =

OpenStudy (anonymous):

-42

Directrix (directrix):

Yes

Directrix (directrix):

What is c^1 times c^6 =

OpenStudy (anonymous):

7

Directrix (directrix):

Not just 7 but what letter raised to the 7th power?

OpenStudy (anonymous):

oh so it c^7

Directrix (directrix):

Correct.

Directrix (directrix):

The final answer is what? Put the c^7 with the -42

OpenStudy (anonymous):

-42c^7

Directrix (directrix):

Very good. Correct.

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

can you tonder me on math